OpenStudy (anonymous):

Prove: sin^4x – cos^4x = sin^2x – cos^2x Step by step please. prove it as simply as you can :)

Parth (parthkohli):

\[\sin^4(x) - \cos^4(x) = \left(\sin^2(x)\right)^2 - \left(\cos^2(x)\right)^2\]Now use an identityt.

Parth (parthkohli):

Identity*

Parth (parthkohli):

\[x^2 - y^2 = (x + y)(x - y)\]

Parth (parthkohli):

Things will begin to unfold.

OpenStudy (anonymous):

\[\Large \sin^4x-\cos^4x\] \[\Large (\sin^2x)^2-(\cos^2x)^2\] you would have to know difference of two squares: \[a^2-b^2=(a+b)(a-b)\] \[\Large (\sin^2x+\cos^2x)(\sin^2x-\cos^x)\] you would have to know that: \[\sin^2x+\cos^2x=1\] \[\Large 1(\sin^2x-\cos^2x)~~=~~\sin^2x-\cos^2x\]

OpenStudy (anonymous):

i meant to write cos^2x in my second step

OpenStudy (anonymous):

Thank you :)

Parth (parthkohli):

Isn't that exactly what I was proving too?